Status Report

ISS On-Orbit Status 17 May 2002

By SpaceRef Editor
May 17, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below.  Day 162 on orbit for Expedition 4.

Elektron update:  After some more troubleshooting
this morning, the Elektron oxygen generator was powered up at
11am-12pm EDT and is running at 32 A (medium setting as indicated for
three persons).  MCC-M has not "declared victory" as
[Leading theory of the Elektron failures has focused on
malfunctioning limit-switch sensors. These redundant microswitches,
activated by the liquid/gas membrane of the water reservoir, indicate
its depletion to the BSSK (signal and command matching unit) before
it can command the inlet valve of the tank open to allow more water
to enter for electrolysis.  It is believed that the contacts
have weakened. A new software algorithm was tested at RSC-Energia’s
NKO test stand today, to be then uplinked by MCC-M to the SM TVM
(terminal computer system). It is not clear at this time whether
today’s restart of the Elektron was accomplished with or without the
new algorithm, which would (1) mask the sensors so that the BSSK does
not see their "tank empty" signal, and (2) emulate the
signal by replacing it with a message after the appropriate time
interval for tank depletion (which depends on selected Elektron ops
mode, in the current case on the 32-amp mode). The algorithm would
henceforth be started every time the Elektron is activated. A second
failure theory has postulated a deformed pressure regulator membrane,
as has been encountered before on Mir and on the ground. The Go for
troubleshooting this possibility has not been given yet.]

The crew has used two SFOGs (solid-fuel oxygen generator) candles
today and two yesterday, leaving a total of 142 on board. Oxygen
partial pressure (ppO2) remains nominal [now measured by the
CSA-CP (compound specific analyzer/combustion products) vs. the SM
gas analyzer (GA)]
. If necessary, SFOGs would be used until 5/22,
after which O2 would be drawn from the Airlock supplies.

Besides the Elektron troubleshooting, CDR Yuri Onufrienko’s main work
today was in the FGB, to take care of the periodic preventive
maintenance of its ventilation system, similar to what he had done in
the SM on 5/8: filter replacements for the redundant systems,
followed by a thorough cleaning of fan screens and flexible air

FE-2 Dan Bursch continued his dedicated work on the BPS (biomass
production system).  The payload is operating nominally, and
Bursch performed a gas calibration and sampling on PGC1 (plant growth
chamber #1), harvesting five samples.  Refilling (priming) of
HCS (humidity control system) reservoir and transfer pump did not
prove necessary, and there is a good possibility that no further
systems priming will have to be done.

The ARIS-ICE (active rack isolation
system/isolation characterization experiment) team is supporting POC
(Payload Operations Center) flight controllers as they control ARIS
from Huntsville for the first time. The tests represent one of
the last ARIS-ICE mission objectives: the handover of the control of the ARIS system to the
POC.  As the test continues this week, the crew will be
contacted periodically since they may notice rack movement.

Onufrienko was provided with instructions for repairing his Nikon-D1X
photo camera which appears to have a problem with the connector pins
for its CompactFlash card units.

A new task dealing with the TVIS (treadmill with vibration
isolation and stabilization) has been added to the crew’s "job
jar" task list.  Ground engineers request video
documentation and audio description of the failed SLD (subject load
device) on the right side.  The visual and audio
characterization of the cable during the pulling actions is expected
to provide valuable insight into the nature of the SLD failure.
 The video tape is to be returned on UF-2.  The crew was
also asked to verify the correct configuration of the eyebolts for
the backup bungee configuration, which were originally installed on
Expedition 3.

Regular routine servicing tasks were performed today by Carl Walz
(IMS delta file preparation, IV-CPDS [intravehicular charged particle
directional spectrometer] power cycling, and autonomous payloads
status checks) and Yuri Onufrienko (SOSH life support systems
inspection, maintenance closure of the Vozdukh vacuum valves, and
checkout of the DC-1 docking compartment’s circuit breaker switch
panel (BVPSO) as well as the state of the fuses in the DC-1 fuse box
panel (BPP).

Onufrienko also removed and replaced an air flow adjustment
damper/door in sleep quarters #1 of the SM.

FE-1 Carl Walz conducted the periodic operability inspection of the
PFEs and PBAs (portable fire extinguishers, portable breathing
apparatus) in both the Lab and the Node.

The two flight engineers completed their FFQ (food frequency
questionnaire) by entering data on food and fluid intake on the MEC
(medical equipment computer) for downlink to the ground at a later

MCC-Moscow continued the ground-commanded tests (sessions 3 and 4) of
the new SA325 set brought up by Progress 7P and installed in the
Regul-OS system by Onufrienko on 5/15.  With it, the Regul,
which handles two-way voice communication, digital command/program
information and telemetry transmission via RGS (Russian ground
stations), now has three redundant sets.

Last night’s attitude change from LVLH to XPOP using the U.S. CMGs
(control moment gyros) was executed with great success. Total
propellants used by the Russian thrusters (for momentum desaturation)
amounted to 3.5 kg.  This is a prop saving of 30% since the same
maneuver performed by SM  thrusters alone was stated by MCC-M to
cost 5 kg.  In addition, the use of the CMGs considerably
simplifies operations by not requiring two "handovers"
between the U.S. and Russian segments.  RSC-Energia agrees that
using the CMGs for this maneuver in the future should be preferred
because of its greater simplicity.  [However, saving
propellants over the long term as a goal becomes increasingly less
desirable because the chemical properties of UDMH (unsymmetrical
dimethyl hydrazine) and NTO (nitrogen tetroxide) deteriorate slowly
with time. Russian ground rule apparently is to use these props
before the end of three years of storage, but even in this time their
combustion process yields decreasing temperatures, i.e., less

At 5/24, MCC-H plans to conduct frequency tests of the S-band
system at the White Sands TDRS ground stations to improve the lock on
the S-band signal from the ISS.  The necessary adjustment in
accumulated frequency drift amounts to 600 Hertz over one year, which
is as expected.

MCC-H performed data collection on the ATU1 (audio terminal unit) in
the Lab yesterday to isolate a problem (loop back error) seen last
year on 12/12.  The ATU continues to work nominally.

MCC-H prepares to swap the roles of the primary and backup GNC MDMs
(guidance, navigation and control computers), making GNC MDM2 the
prime machine, prior to UF-2.  [Rationale:  the GNC1,
currently Primary, and the Primary C&C MDM3 (command and control
computer #3) are on the same power feed, while GNC2 is on a separate
channel. This arrangement  would deactivate both of the
Primaries in the event of a power failure, and the backup C&C
would then invoke an attitude control handover to the Russian
segment. If this occurred after Shuttle docking, the SM motion
control system would begin reacting against the Orbiter’s attitude
control systems.]

An RPC (remote power controller) for Node lighting tripped and
was restored to nominal function after data collection. Cause of the
"blown fuse" is unknown at this time (no current spike
indicated), but the switch has tripped once before.

CSA (Canadian Space Agency) reports readiness of the new WR (wrist
roll) joint for the joint replacement "surgery" planned on
the SSRMS for UF-2 EVA-3.  But work is still going on to update
the JCS (joint control software) after it was discovered that it
contains four software files with data on unique physical
characteristics of the joint (such as motor specifics, e.g.
resistance, induction, etc., and offset value of the particular
joint’s angle resolver).  These parameters will be changed and
the changes tested, with uplink date to be determined.

Today’s target areas for the U.S. CEO (crew earth observation)
program were W. Mediterranean Dust and Smog (a large
weather system over the NE Atlantic has set up northerly circulation,
and Moroccan dust is being carried out over the western
Mediterranean. Crew was to look right, toward the Strait of
Gibraltar, as they exited the coast of Spain to capture the North
African coast and blowing dust)
, Angolan Biomass Burning
(oblique views to right of track, as ISS reached the coast just
south of Luanda, to document smoke moving offshore from the interior
of Angola)
, Industrialized Southeastern Africa (the
Johannesburg-Pretoria urban center was just left of track. Crew was
asked to look right, down the broad valley of the Orange and Vaal
rivers to record any smog moving south toward the Karroo Ranges)
Cuban Fires, Smoke (this pass took the station over the
western end of Cuba. While ISS attitude might have precluded
photographing the persistent fires there, the crew should have been
able to record Cuban smoke movement over the Gulf of Mexico. In
addition, they were to look right and document the current
distribution of Central American smoke. Clouds appear to be moving
into Yucatan, and there’s hope that rains might quench the
, Lakes, E. Sierra Nevada Watershed (context
photos, particularly of Mono Lake and the Salton Sea, were needed;
both to left of track. Snowmelt from the Sierra Nevada is one of the
water sources for these lakes; crew was to document the extent of
snow in the Sierras, in the wake of recent storms)
CEO images can be viewed at the website

U.S. and Russian
Segment Status
(as of 1:30 pm EDT):

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is On (intermittently?); SFOG
candles (TGKs) are being used.  Vozdukh CO2 scrubber is ON in
MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA CO2
scrubber is Off.  BMP Harmful Impurities unit: Absorbent bed #1
in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 756,
temperature (deg C) — 28.1, ppO2 (mmHg) — 146.1, ppCO2 (mmHg) —
2.5 (probably invalid)
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 21.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 23.0.
Node:  Pressure (mmHg) — 747.07, temperature (deg C) —
24.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 749.26, temperature (deg C)
— 24.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 749.16,
temperature (deg C) — 24.3; shell heater temp (deg C) — 24.0, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 23.4
PMA-2:  Shell heater temp (deg C) — 12.9

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: Battery #6 is offline (failed); all other batteries (7)
are in “partial charge” mode.
FGB: Battery #3 is disconnected; all other batteries (5) are in
“partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 at Progress viewing position, with Keep Alive
power on both strings.
RWS (robotics workstations) are Off.
SSRMS Prime string Wrist Roll (WR) joint to be replaced on UF-2.

ISS Orbit (as of this morning, 7:22 am EDT [=
Mean altitude — 389.3 km
Apogee — 393.8 km
Perigee — 385.8 km
Period — 92.3 min.
Inclination (to Equator) — 51.63 deg
Eccentricity — 0.0006646
Orbits per 24-hr. day — 15.59
Altitude decrease — 300 m (mean) in last 24 hours
Solar Beta Angle — 32.8 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 19923
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.6 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.